the measure of the supplement of an angle is 25degress more than 7 times the measure of the angle.To the nearest hundredth. what is the measure of the angle?
Found 2 solutions by rfer, ewatrrr:
Answer by rfer(16322)

x+7x+25=180
8x=155
x=19.38
7x+25=160.62

Answer by ewatrrr(24785)


Hi
Let x represent "the angle" Supllement would be (180-x)
(180-x) = 7x + 25
180 - x = 7x + 25
155 = 8x
19.38dgrees = x